OpenStudy (anonymous):

What is the prime factorization of 54? A. 9 • 6 B. 22 • 32 C. 33 • 2 D. 32 • 6

OpenStudy (charlotte123):

divide 54 by 3. = 3*18 divide the 18 by 2 = 9*2 divide the 9 by 3 =3*3 so your answer is 2*3*3*3 --> 2 3^3 (or in this case, C)

OpenStudy (charlotte123):

thank u!